# Turbulence in stratified rotating topographic wakes

### v1

@@ Abstract
Turbulence generation mechanisms in stratified, rotating flows past three-dimensional (3D) topography remain underexplored, particularly in submesoscale (SMS) regimes critical to geophysical applications. Using turbulence-resolving large-eddy simulations, we systematically dissect the interplay of stratification and rotation in governing the dynamics of wake turbulence. Our parametric study reveals that turbulent dissipation in the near wake is dominated by two distinct instabilities: (1) vertical shear-driven Kelvin-Helmholtz instability (KHI), amplified by oblique dislocation of Kármán vortices under strong stratification, and (2) centrifugal/inertial instability (CI), which peaks at intermediate rotation rates (Rossby number order unity, SMS regime) and relatively weaker stratification. Notably, strong rotation dampens vertical shear and weakens KHI-driven turbulence, while strong stratification imposes smaller vertical length scales that restricts CI-driven turbulence. By quantifying dissipation across a broad parameter space of stratification and rotation, predictive relationships between the environmental parameters and instability dominance is established. These findings highlight the regime dependence of instability mechanisms and may inform targeted observational campaigns and numerical models of oceanic and atmospheric wakes.
@@ 1 Introduction
[…]
@@
Motivated by the need to understand and model topographic wake turbulence, we employ a suite of high-resolution LES with the objective being the characterization of turbulence and mixing. The physical model is a three-dimensional (3D) conical topography sitting on the ocean bottom, as shown in figure 1 of Liu et al. (2024). The topography has a base diameter $D=150\,{\rm m}$, height $h=150\,{\rm m}$ and a slope approximately $30^{\circ}$. This setting represents a model of a steep topography, such as seamounts. The fluid is linearly stratified and the coordinate system rotates at a constant rate ($f$-plane). The flow impinging on the obstacle is a steady current and the focus of this work will be the comprehensive effects of stratification, rotation and 3D topography, and how they affect the vortex dynamics and turbulent dissipation. Such simplifications allow a focus on wake turbulence without the loss of generality.
@@ 2.1 Large-eddy simulations
the buoyancy force is $b=-{\rho^{*}g}/{\rho_{0}}$, and the Coriolis force is $\bm{f}_{c}=f_{c}{\bm{e}}_{z}$ ($f_{c}$ is a constant; $f$-plane). Here $p^{*}$ and $\rho^{*}$ are the deviations from the hydrostatic and geostrophic balances, and $\nu$ and $\kappa$ are the molecular viscosity and diffusivity. The subscript ‘sgs’ denotes the sub-grid scale contributions to the momentum and scalar transport, and are modeled with the WALE (wall-adapted local eddy-viscosity; Nicoud and Ducros (1999)) method in an LES approach.
@@
$Fr=\frac{U_{\infty}}{Nh},\,Ro=\frac{U_{\infty}}{f_{c}D},\,Re=\frac{U_{\infty}D}{\nu},$ (5)
are the key controlling parameters that will be systematically varied in this study. We opt to use a Southern Hemisphere rotation and $f_{c}$ represents the absolute value of a negative Coriolis frequency.
@@
Table 1: Parameters of the computational study which has 4 series corresponding to $Fr=0.075$, 0.15, 0.30 and 0.40 with varying $Ro$ within a series. Each series has a fixed combination of $Fr$ and $Re$ such that the stratification scale based Reynolds number, $Re_{N}=U_{\infty}^{2}/\nu N=FrRe(h/D)$, is a constant at $Re_{N}=900$. Dimensional reference values for $N,f_{c}$ and $U_{\infty}$ are given for cases with $1\leq Bu\leq 400$. The ranges of the buoyancy frequency and the Coriolis frequency are $0.5\times 10^{-3}\,{\rm s^{-1}}\leq N\leq 0.5\times 10^{-2}\,{\rm s^{-1}}$ and $2.5\times 10^{-5}\leq f_{c}\leq 1.4\times 10^{-4}$ (latitudes between 10 and 75). The hill has base diameter $D=500$ m and height $h=150$ m.
Figure 1: Parameter space spanned by $(Fr,Ro)$. The horizontal axis is the inverse Froude number, $Fr^{-1}$, and the vertical axis is the Rossby number, $Ro$. Dashed lines represent constant Burger numbers, $Bu=(Ro/Fr)^{2}=1,4,25,100,400$, increasing from the lower left to the upper right. For each $Fr$, the four cases enclosed by $1\leq Bu\leq 400$, will be analyzed in detail.
2.2 Parameter selection
In order to cover a wide range of combinations of stratification and rotation, a parameter sweep is conducted in the $Fr$$Ro$ space, as shown in Table 1 and Fig. 1. The non-dimensional parameters are selected as follows.
Although the stratification in the deep ocean has seasonal and geographical variability, a buoyancy frequency of $N=0.5\times 10^{-3}\,{\rm s^{-1}}$ can be taken as representative. A mountain height of 1 km and a current speed between $0.2\,{\rm m\,s^{-1}}$ and $0.02\,{\rm m\,s^{-1}}$ , leads to $Fr$ between 0.4 and 0.04. Accordingly, four values of $Fr$, from 0.075 to 0.40, that span half an order of magnitude, are selected. All cases lie in the flow-around regime (Hunt and Snyder, 1980; Chomaz et al., 1993), and they represent moderate to relatively strong stratification. The focus is on steep underwater topography that are known sites of elevated mixing and, accordingly, $h/D=0.3$.
@@ 2.2 Parameter selection
[…]
@@
There is an additional (but not independent) non-dimensional number, the Burger number ($Bu=(Ro/Fr)^{2}=(Nh/f_{c}D)^{2}$), commonly used to describe the relative importance of rotation to stratification in geophysical flows. When $Bu\leq O(1)$, rotation dominates and when $Bu>O(10)$, the effect of rotation is relatively small compared to the dominance of stratification. The range of Burger number of $1\leq Bu\leq 400$, shown by the four diagonal dashed lines (each has a constant value of $Bu$) in Fig. 1, will be examined in detail.
@@
The Reynolds number simulated should be sufficiently high so that the flow becomes turbulent and as many instabilities are triggered as possible. At $Fr=O(0.1)$, the regime studied here, the wake is in the flow-around vortex shedding regime, where the horizontal components dominate the turbulent kinetic energy (TKE). However, there are vertical structures and (oblique) dislocations (see Fig. 2) of these horizontal motions, leading to strong vertical shear and the potential for KHI. This scenario is a generic feature of vertically stratified horizontal shear flows (Billant and Chomaz, 2000; Basak and Sarkar, 2006) and such a turbulence generation mechanism is referred to as a shortcut in the transition (Deloncle et al., 2008; Waite and Smolarkiewicz, 2008), since the flow is originally sheared in the horizontal directions.
@@
The normal-mode KHI requires at least one point in the flow to have a gradient Richardson number less than 1/4 (Miles, 1961; Howard, 1961). Using the viscous length scale between dislocated layers measured by Basak and Sarkar (2006), approximately $l_{d}=15\sqrt{\nu/N}$, and a velocity difference $2U_{\infty}$ between oppositely flapping vortex shedding, the resulting maximum vertical shear is $S_{v}=2U_{\infty}/l_{d}$. In order to reach a marginal gradient Richardson number, $Ri_{g}=N^{2}/S_{v}^{2}=1/4$, a minimum Reynolds number of $Re_{N}=U_{\infty}^{2}/\nu N=225$ is required.
[…]
@@ 3 Dissipation features and their link to flow instabilities
Topographic wakes are enriched by multi-scale and multi-physics interactions. Eddies whose sizes range from planetary scale, for example, MS of $O(100{\rm km})$ to SMS of $O(0.1$${\rm km})$ to turbulent eddies that are only a few centimeters in size evolve concurrently in the near and far wakes of the topography. The flow exhibits a Kármán shedding pattern in horizontal cross-sections with some vertical structures (Liu et al., 2024). In this work, we are concerned with wake turbulence, where there are several physical mechanisms that can potentially lead to the destabilization of the flow and the breakdown into turbulence in the wake. In the vertical direction, intense shear (compared to the strength of stratification) can lead to KHI, while the source of shear varies from the dislocated vortex shedding at different elevations to the undulating lee wave near the apex of the obstacle. For the horizontal motions, anticyclonic vortices and shear are both subject to CI. These two key mechanisms are determined to be operative in wake turbulence later on and their parametric dependence is investigated.
@@ 3.1 Turbulence statistics and quantification
Definitions and notations that are used in the statistical analysis are as follows. An instantaneous signal is decomposed into time-averaged and fluctuation parts as
@@
[…]
@@ 3.2 Contribution of the vertical shear instability (KHI)
(a) $Fr=0.075,Ro=0.75$
(b) $Fr=0.15,Ro=0.75$
(c) $Fr=0.30,Ro=0.75$
(d) $Fr=0.40,Ro=0.75$
Figure 3: Instantaneous dissipation rate $\varepsilon$ (colored) and its time-averaged $\langle{\varepsilon}\rangle$ (TKE dissipation, gray), probed on the line $y=0,x=1$. Time instances are the at the same as those in Fig. 2(a-d).
@@
Here $u,v$ and $b$ are instantaneous quantities. A value of $Ri_{g}=1/4$ typically indicates a marginal instability state due to vertical shear. It is clear that the shear-unstable regions coincide with the strongest dissipation, suggesting that the KHI is active.
@@
Comparing Figs. 2(b,e,f), at $Fr=0.15$ and various $Ro$, it can be seen that the SMS cases $Ro=0.75,1.5$ are similar while the MS case $Ro=0.15$ shows fewer large-$\varepsilon$ patches. At $Ro=0.75,1.5$, layers of tilted coherent vortex structures are shed and the shear instability serves as the main contributor to turbulence. When rotation is strong ($Ro=O(0.1)$), vertical gradients are significantly reduced and columnar vortices emerge which, further downstream, advect as stratified Taylor columns (Liu et al., 2024). Figure 2(e) shows turbulent dissipation associated with these columnar vortices, which is significantly weaker than the dissipation associated with the slanted layers, shown in (b,f). The contrast indicates that the form of the coherent structures, in turn, influenced by rotation, can also influence turbulence intensity. This might be regarded as the indirect effect of rotation on turbulence.
Another contributor to turbulence is the hydraulic jet that plunges below the mountain crest and also sets up a near-field internal wave response. The topmost portion of the mountain, from its crest to $U_{f}/N$ below (Winters and Armi, 2012), participates in the jet and the lee wave, whose dominant vertical wavelength is $2\pi U_{m}/N$ (Klymak et al., 2010). Here $U_{f}$ and $U_{m}$ are the freestream and mean velocity, respectively. The undulating jet and the adjacent wave region are found to be a dissipation hotspot that becomes more pronounced with increasing $Ro$ and $Fr$ of the examined cases. The flow in the hotspots breaks down through the KHI and forms rolled-up billows that are evident in the insets of Figs. 2(g-h), corresponding to Fr030Ro3 and Fr040Ro4.
Moreover, in all cases, there is turbulent dissipation due to the unsteady internal waves propagating into the background, as shown by the yellowish colors above the top of the hill, albeit smaller than the dissipation in the wake (note the logarithmic color scale in Fig. 2). While the wave dissipation represents a different process that is also important, wake dissipation will be the focus of this work. In the vertical center plane, TPE dissipation (not shown) has a spatial distribution that aligns well with TKE dissipation. Accompanying movies of $\varepsilon$ for vertical center planes as shown in panels (a-d) can be found in the supplementary materials.
Figure 4: TKE dissipation $\langle\varepsilon\rangle$ measured at $y=0,\,x=1$. (a) $Fr=0.075$, (b) $Fr=0.15$, (c) $Fr=0.30$, and (d) $Fr=0.40$.
@@
Figure 5: Contours of the potential vorticity, ${\it\Pi}/N^{2}f_{c}$, in the horizontal plane at $z/h=0.25$. Panels (a-d): $Fr=0.075,0.15,0.30,0.40$, respectively, and $Ro=0.75$ for all panels. The background value is unity for all panels, corresponding to normalized background PV (unity) due to the system rotation.
@@
Figure 4 compares the dissipation profiles at $x/D=1,y=0$ for different $(Fr,Ro)$ cases. The dual roles of rotation can be seen. On one hand, as shown in Fig. 4(c), case Fr030Ro075 with rotation at SMS has the largest dissipation at the centerline relative to $Ro$ values that are higher or lower. The peak at SMS Ro, which will also be shown in volume-integrated $\varepsilon$ later, is linked to CI. On the other hand, in Fig. 4(a-b), very strong rotation ($Ro=0.15$) significantly reduces $\varepsilon$ by almost an order of magnitude compared to $Ro\geq O(1)$, through the reduction of the vertical velocity gradients and the consequent vertical shear instability. Moving from Fig. 4(b) to (c), the effect of rotation appears to be non-monotonic due to the reasons mentioned above, but a more comprehensive measure of the dissipation in the 3D domain is required and the examination of the CI instabilities is needed. These investigations are reserved for the next sections.

### v2

@@ Abstract
Turbulence generation mechanisms in stratified, rotating flows past three-dimensional (3D) topography remain under-explored, particularly in submesoscale (SMS) regimes critical to geophysical applications. Using turbulence-resolving large-eddy simulations, we systematically dissect the interplay of stratification and rotation in governing the dynamics of wake turbulence. Our parametric study reveals that turbulent dissipation in the near wake is dominated by two distinct instabilities: (1) vertical shear-driven Kelvin-Helmholtz instability (KHI), amplified by oblique dislocation of Kármán vortices under strong stratification, and (2) centrifugal/inertial instability (CI), which peaks at intermediate rotation rates (Rossby number order unity, SMS regime) and relatively weaker stratification. Notably, strong rotation dampens vertical shear and weakens KHI-driven turbulence, while strong stratification imposes smaller vertical length scales that restrict CI-driven turbulence. By quantifying dissipation across a broad parameter space of stratification and rotation, predictive relationships between the environmental parameters and instability dominance are established. These findings highlight the regime dependence of instability mechanisms and may inform targeted observational campaigns and numerical models of oceanic and atmospheric wakes.
@@ 1 Introduction
Recent studies (Johnston et al., 2019; MacKinnon et al., 2019; Rudnick et al., 2019) during the Flow Encountering Abrupt Topography (FLEAT) initiative have shed light on the eddy field surrounding Palau Island which lies in the path of the North Equatorial Current (NEC). Cyclonic and anti-cyclonic eddies spanning a wide range of Rossby numbers ($Ro\approx 0.3$$30$) are observed in the lee of the island with diameters ranging from 1-10 km (submesoscale, SMS) to several hundred km (mesoscale, MS), which are comparable to the width of the island.
@@
Motivated by the need to understand and model topographic wake turbulence, we employ a suite of high-resolution LES with the objective being the characterization of turbulence and mixing. The physical model is a three-dimensional (3D) conical topography sitting on the ocean bottom, as shown in figure 1 of Liu et al. (2024). The topography has a base diameter $D=500\,{\rm m}$, height $h=150\,{\rm m}$ and a slope approximately $30^{\circ}$. This setting represents a model of a steep topography, such as seamounts. The fluid is linearly stratified and the coordinate system rotates at a constant rate ($f$-plane). The flow impinging on the obstacle is a steady current and the focus of this work is on the comprehensive effects of stratification, rotation and 3D topography, and how they affect the vortex dynamics and turbulent dissipation. Such simplifications allow a focus on wake turbulence without the loss of generality.
@@ 2.1 Large-eddy simulations
the buoyancy is $b=-{\rho^{*}g}/{\rho_{0}}$. Here $p^{*}$ and $\rho^{*}$ are the deviations from the hydrostatic and geostrophic balances, and $\nu$ and $\kappa$ are the molecular viscosity and diffusivity. The subscript ‘sgs’ denotes the sub-grid scale contributions to the momentum and scalar transport, and are modeled with the WALE (wall-adapted local eddy-viscosity; Nicoud and Ducros (1999)) method in an LES approach. The Coriolis force involves the planetary rotation vector, $\bm{f}_{c}=f_{c}{\bm{e}}_{z}$, where the sign of the Coriolis frequency $f_{c}$ is negative (Southern Hemisphere), and the magnitude of $f_{c}$ is a constant as per the $f$-plane approximation, in the present work. We opt to use a Southern Hemisphere rotation but the results are equally applicable to both hemispheres with the distinction between cyclones and anticyclones carefully made.
@@
$Fr=\frac{U_{\infty}}{Nh},\,Ro=\frac{U_{\infty}}{|f_{c}|D},\,Re=\frac{U_{\infty}D}{\nu},$ (5)
are the key controlling parameters that will be systematically varied in this study, where the definition of $Ro$ involves only the magnitude of $f_{c}$.
@@
Figure 1: Topographic wake generated by a conical seamount. Red and blue contours show the isosurfaces of vertical vorticity, $\omega_{z}D/U_{\infty}=\pm 1$, representative of oblique vortex shedding structures. The sidewall projection shows the turbulent dissipation rate in logarithmic scale (for the colorbar see Fig. 3), at the same time instance and on the central plane ($y=0$). The upper panels show specific regions of the flow, featuring distinct physical mechanisms. The flow goes from left to right. Case Fr030Ro3.
2.2 Flow features and parameter selection
Topographic wakes are enriched by multi-physics and multi-scale interactions. Fig. 1 depicts the flow patterns. As the flow passes the obstacle, a hydraulic response is established near the top of the seamount resulting in downslope flow and an undulatory jet, both with strong vertical shear. Moreover, internal waves are transmitted into the stratified medium. In the wake, eddies of different scales ranging from large-scale coherent wake eddies to dissipative eddies coexist and interact. At the large scale, the wake exhibits Kármán shedding patterns at different heights. It is intuitive to think that, in strongly stratified flows, Kármán shedding is vertically decoupled into independent, stacked layers. However, the vortex shedding structures were found to share a vertically independent global frequency in the $Fr=0.15$ seamount wakes investigated by Liu et al. (2024). When rotation is not strong, the coherent structures are oblique ‘surfboards’. They lean into the streamwise direction and form a shallow angle to the horizon, as shown in Fig. 1. On one hand, the existence of these coherent structures is robust to the presence of near wake turbulence. On the other hand, the vortex and shear structures associated with them provide a breeding bed for various instabilities.
In this work, we are mainly concerned with wake turbulence, where several physical mechanisms can potentially lead to the destabilization of the wake and the breakdown into turbulence. Vertical shear that is sufficiently strong compared to $N$ will be unstable to KHI, where the source of shear can be in between opposite-signed ‘surfboard’ vortices or from the intensified shear of the hydraulic jet (HJ). In the horizontal directions, anticyclonic vortices and shear layers are both subject to CI. These two key mechanisms – KHI and CI – are determined to be operative in wake turbulence later on and their parametric dependence is investigated.
@@ 2.2 Flow features and parameter selection
Table 1: Parameters of the computational study which has 4 series corresponding to $Fr=0.075$, 0.15, 0.30 and 0.40 with RoRo varying within a series. Each series has a fixed combination of $Fr$ and $Re$ such that the stratification scale based Reynolds number, $Re_{N}=U_{\infty}^{2}/\nu N=FrRe(h/D)$, is a constant at $Re_{N}=900$. Dimensional reference values for $N,|f_{c}|$ and $U_{\infty}$ are given for buoyancy frequency in the range $0.5\times 10^{-3}\,{\rm s^{-1}}\leq N\leq 0.5\times 10^{-2}\,{\rm s^{-1}}$ and the Coriolis frequency in the range $2.5\times 10^{-5}\leq|f_{c}|\leq 1.4\times 10^{-4}$ (latitudes between 10 and 75), based on $D=500$ m and $h=150$ m. Cases with reference values not provided (marked as N/A) are still dynamically relevant with different values of $D$ and $h$ but do not have reasonable dimensional values for $D=500$ m and $h=150$ m.
Figure 2: Parameter space spanned by $(Fr,Ro)$. The horizontal axis is the inverse Froude number, $Fr^{-1}$, and the vertical axis is the Rossby number, RoRo. Dashed lines represent constant Burger numbers, $Bu=(Ro/Fr)^{2}=1,4,25,100,400$, increasing from the lower left to the upper right. For each $Fr$, the four cases enclosed by $1\leq Bu\leq 400$, will be analyzed in detail.
In order to cover a wide range of combinations of stratification and rotation, a parameter sweep is conducted in the FrFrRoRo space, as shown in Table 1 and Fig. 2. The non-dimensional parameters are selected as follows.
Although the stratification in the deep ocean has seasonal and geographical variability, a buoyancy frequency of $N=0.5\times 10^{-3}\,{\rm s^{-1}}$ can be taken as representative. A mountain height of 1 km and a current speed between $0.2\,{\rm m\,s^{-1}}$ and $0.02\,{\rm m\,s^{-1}}$ leads to $Fr$ between 0.4 and 0.04. Accordingly, four values of $Fr$, from 0.075 to 0.40, that span half an order of magnitude, are selected. All cases lie in the flow-around regime (Hunt and Snyder, 1980; Chomaz et al., 1993), and they represent moderate to relatively strong stratification. The focus is on steep underwater topography that are known sites of elevated mixing and, accordingly, a representative value of $h/D=0.3$ corresponding to a slope of approximately 30 is chosen. Here, ‘steep’ refers to the following dynamical aspects: (1) the terrain slope exceeds typical angles of oceanic internal wave propagation, and (2) a substantial part of the wake is in the flow-around, vortex-shedding regime.
@@
The FrFrRoRo parameter space also extends the scope of previous work employing ROMS. The work of Perfect et al. (2020) focused on large topographies in the MS with $Fr,Ro\sim O(0.01-0.1)$. Similarly, Srinivasan et al. (2021) studied strong rotation and strong stratification ($Fr=0.02,\,0.025\leq Ro\leq 1$, in the present definition). In the limits of strongly rotating and/or stratified flows, the hydrostatic assumption in ROMS generally works well for the large-scale motions and both stratification and rotation act as stabilizing factors for smaller-scale motions (Perfect et al., 2020). When hydrodynamic instabilities and turbulence are of interest as is the case here, non-hydrostatic simulations are required to resolve them and, thus, elucidate their qualitative and quantitative dependence on $(Fr,Ro)$.
@@
There is an additional (but not independent) non-dimensional number, the Burger number ($Bu=(Ro/Fr)^{2}=(Nh/f_{c}D)^{2}$), commonly used to describe the relative importance of rotation to stratification in geophysical flows. When $Bu\leq O(1)$, rotation dominates and when $Bu>O(10)$, the effect of rotation is relatively small compared to the dominance of stratification. The range of $1\leq Bu\leq 400$, shown by the four diagonal dashed lines (each has a constant value of $Bu$) in Fig. 2, will be examined in detail.
@@
The chosen Reynolds number should be sufficiently high so that the flow becomes turbulent and as many instabilities are triggered as possible. At the same time, $Re$ cannot be too high so as to avoid numerical dispersion errors arising from flow instabilities that are poorly resolved on the chosen grid. At $Fr=O(0.1)$, the regime studied here, the wake is in the flow-around vortex shedding regime, where the horizontal components dominate the turbulent kinetic energy (TKE). However, there are vertical structures and (oblique) dislocations (see Fig. 3) of these horizontal motions, leading to strong vertical shear and the potential for KHI. This scenario is a generic feature of vertically stratified horizontal shear flows (Billant and Chomaz, 2000; Basak and Sarkar, 2006) and such a turbulence generation mechanism is referred to as a shortcut in the transition (Deloncle et al., 2008; Waite and Smolarkiewicz, 2008), since the flow is originally sheared in the horizontal directions.
@@
The normal-mode KHI requires at least one point in the flow to have a gradient Richardson number less than 1/4 (Miles, 1961; Howard, 1961). From Fig. 1, it can be seen that a surfboard structure lies on top of another with oppositely-signed vorticity (near $x/D=4$). This scenario constitutes an oblique dislocation. When the direction of spanwise wake flapping, indicated by the sign of Kármán vortices, is opposite across a vertical distance, intense vertical shear is generated. Using the viscous length scale between dislocated layers measured in Basak and Sarkar (2006), approximately $l_{d}=15\sqrt{\nu/N}$, and a velocity difference $2U_{\infty}$ between oppositely flapping wake vortices separated by this distance, the resulting vertical shear is $S_{v}=2U_{\infty}/l_{d}$. The Reynolds number at which this shear is marginally stable, such that $Ri_{g}=N^{2}/S_{v}^{2}=1/4$, is $Re_{N}=U_{\infty}^{2}/\nu N=225$.
Thus, $Re_{N}=U_{\infty}^{2}/\nu N=FrRe(h/D)$, instead of $Re$ or $Fr$ alone, serves as the a priori indicator of transition to turbulence through KHI in the present wakes. It can also be interpreted as the Reynolds number of turbulent eddies whose size does not exceed the stratification length scale $\sim U_{\infty}/N$. Such an estimate has the following implications. First, the Reynolds number required to have turbulence increases with stratification and can be quite large at strong stratification. For example, for $Fr=0.15$, the constraint of $Re_{N}>225$ requires a molecular Reynolds number of $Re>5000$, which is higher than what many laboratory experiments and numerical simulations of the large-eddy and direct numerical simulation (LES and DNS) classes have reached. Second, if a similar dynamic range of turbulence is desired for a series of experiments, one also needs to consider matching a lower $Fr$ with a higher $Re$ and vice versa so that $Re_{N}$ is comparable, in addition to fixing one and varying another. We base the selection of the Reynolds number on such an estimate and hold $Re_{N}=900$ constant across all cases as an a priori choice of a buoyancy-based Reynolds number. The value of $Re_{N}=900$ is four times the estimated critical value so as to ensure fully-developed turbulence. It will be shown a posteriori that KHI is active in all cases. The highest Reynolds number is up to Re=40  [^0] Re=40\,000 at $Fr=0.075$. A Reynolds number sensitivity study is provided in Appendix B to show that turbulent dissipation statistics approach $Re$-independence in the present simulations.
@@ 3.1 Turbulence statistics and quantification
Definitions and notations that are used in the statistical analysis are as follows. An instantaneous signal is decomposed into the mean and fluctuation as
@@
where the bracket $\langle\cdot\rangle$ denotes time average. The ensemble size for the time average is $N_{t}\approx 375$ when snapshots of 2D planes are analyzed, and is $N_{t}\approx 30$ when snapshots of 3D boxes are analyzed. The turbulent kinetic and potential energy (TKE and TPE) are
@@
Figure 4: Instantaneous dissipation rate $\varepsilon$ (colored) and its time-averaged $\langle{\varepsilon}\rangle$ (TKE dissipation, gray), probed on the line $y=0,x/D=1$. Time instances are the same as those in Fig. 3(a-d).
@@ 3.2 Contribution of the vertical shear instability (KHI)
Here $u,v$ and $b$ are instantaneous quantities. A value of $Ri_{g}=1/4$ typically indicates a marginal instability state due to vertical shear. It is clear that the shear-unstable regions coincide with the strongest dissipation, suggesting that KHI is active.
@@
Comparing Figs. 3(b,e,f), at $Fr=0.15$ and various $Ro$, it can be seen that the SMS cases $Ro=0.75,1.5$ are similar while the MS case $Ro=0.15$ shows fewer large-$\varepsilon$ patches. At $Ro=0.75,1.5$, layers of tilted coherent vortex structures are shed and the shear instability serves as the main contributor to turbulence. When rotation is strong ($Ro=O(0.1)$), vertical gradients are significantly reduced and columnar vortices emerge which, further downstream, advect as stratified Taylor columns (Liu et al., 2024). Figure 3(e) shows turbulent dissipation associated with these columnar vortices, which is significantly weaker than the dissipation associated with the slanted layers in (b,f). The contrast indicates that the form of the coherent structures, in turn, influenced by rotation, influences turbulence intensity. This might be regarded as the indirect effect of rotation on turbulence.
Another contributor to turbulence is the HJ that plunges below the mountain crest and also sets up a near-field internal wave response. The topmost portion of the mountain, from its crest to $U_{f}/N$ below (Winters and Armi, 2012), participates in the jet and the lee wave, whose dominant vertical wavelength is $2\pi U_{m}/N$ (Klymak et al., 2010). Here $U_{f}$ and $U_{m}$ are the freestream and mean velocity, respectively. The undulating jet and the adjacent wave region constitute a dissipation hotspot that becomes more pronounced with increasing $Ro$ and $Fr$ of the examined cases. The flow in the hotspots breaks down through KHI and forms rolled-up billows that are evident in the insets of Figs. 3(g-h), corresponding to Fr030Ro3 and Fr040Ro4.
Moreover, in all cases, there is turbulent dissipation due to the unsteady internal waves propagating into the background, as shown by the yellowish colors above the top of the seamount, albeit smaller than the dissipation in the wake (note the logarithmic color scale in Fig. 3). While the wave dissipation represents a different process that is also important, wake dissipation will be the focus of this work. In the vertical center plane, TPE dissipation (not shown) has a spatial distribution that aligns well with TKE dissipation. Accompanying movies of $\varepsilon$ for vertical center planes as shown in panels (a-d) can be found in the supplementary materials.
Figure 5: TKE dissipation $\langle\varepsilon\rangle$ measured at y=0,x/D=1y=0,\,x/D=1. (a) $Fr=0.075$, (b) $Fr=0.15$, (c) $Fr=0.30$, and (d) $Fr=0.40$.
@@
Figure 6: Contours of the normalized potential vorticity, ${\it\Pi}/N^{2}f_{c}$, in the horizontal plane at $z/h=0.25$. Note that $f_{c}$ is signed and $f_{c}<0$ in the present work, such that negative normalized PV resides on the $y<0$ side, in general. Panels (a-d): $Fr=0.075,0.15,0.30,0.40$, respectively, and $Ro=0.75$ for all panels. The background value is unity for all panels, corresponding to normalized background PV (unity) due to the system rotation.
@@
Figure 5 compares the dissipation profiles at $x/D=1,y=0$ for different $(Fr,Ro)$ cases. The dual role of rotation can be seen. On one hand, as shown in Fig. 5(c), case Fr030Ro075 with an intermediate rotation rate has the largest dissipation at the centerline compared to other $Ro$ values that are higher or lower. This is also confirmed later to be the case in volume-integrated $\varepsilon$ and shown to be linked to CI. On the other hand, in Fig. 5(a-b), very strong rotation ($Ro=0.15$) significantly reduces $\varepsilon$ by almost an order of magnitude compared to $Ro\geq O(1)$, through the reduction of the vertical velocity gradients and the consequent vertical shear instability. Moving from Fig. 5(b) to (c), the effect of rotation appears to be non-monotonic due to the reasons mentioned above, but a more comprehensive measure of the dissipation in the 3D domain is required and the examination of the CI is needed. These investigations are reserved for the next sections.
It is shown in Fig. 3(g,h) that the vertical shear that leads to KHI can originate from the vortex structures in the wake and the HJ near the top of the seamount. These two mechanisms are distinct and both play an important role in the dynamics of topographic flows. A quantitative assessment reveals their relative contribution. In the center plane ($y=0$), the HJ region (appropriately 11.2Fr<z/h<11-1.2Fr<z/h<1) contributes 24-32% and 35-51% of the planar dissipation in cases Fr030 and Fr040, respectively. Note that this range of 11.2Fr<z/h<11-1.2Fr<z/h<1 is nearly half of the wake height ($0<z/h<1$) in case Fr040 so the large percentage is not surprising. In volumetric measures (which are analyzed in detail in section 4), the similarly defined HJ region contributes less than 3% to the total volume-integrated dissipation in cases Fr030 and Fr040 and less at lower $Fr$. The implication is two-fold. On one hand, the significance of HJ to dissipation cannot be ignored, especially in the vertical center plane – where it is the strongest. On the other hand, in flows around 3D topographies, its volumetric contribution is much smaller compared to the wake core, unlike in 2D topographic flows.